OpenStudy (anonymous):

one printer can do a printing job in 2hr. Another printer can do the same job in 8hr. How long does it take the two to do the job together?

OpenStudy (anonymous):

\[\frac{8\times 2}{8+2}\] \[\frac{16}{10}=1.6\]

OpenStudy (anonymous):

the answers to choose from are 1&5/11hr 1&3/5hr 2&2/3hr or 1/10hr

OpenStudy (anonymous):

1.6 is one and three fifths

OpenStudy (phi):

As you can see from sat's answer, there is a simple formula to solve these types of problems. Hopefully you can just learn it. But here is a way to do the problem that might help? when they say a printer does a job in 2 hours, you can think of it as 1 job per 2 hours, or 1/2 job per hour. If you multiply by the number of hours it works, you get the number of jobs it does. For example, 1/2 job per hour times 2 hours = 1 job. With 2 printers, working together, where one does 1/2 job per hour, and the other does 1/8 job per hour, we find the number of jobs by adding: 1/2 * t + 1/8 * t = # of jobs Here t is the number of hours they print. Of course, you want just 1 job. so 1/2 t + 1/8 t = 1 can you solve for t? Put them over a common denominator 8: 4/8 t + 1/8 t = 1 5/8 t =1 t = 8/5 hour change 8/5 hour to 1 and 3/5 hour or 1 hour 36 minutes